OpenStudy (anonymous):

How do I solve:? Show that 480 W of power is expended by a weightlifter when lifting a 60-kg barbell a vertical distance of 1.2 m in a time interval of 1.5 s.

OpenStudy (abhisar):

Power = \[\frac{ Work~done }{ Time~taken }\]

OpenStudy (anonymous):

480/1.5?

OpenStudy (abhisar):

\[Work = F X distance\]

OpenStudy (anonymous):

680? 480?

OpenStudy (anonymous):

explain

OpenStudy (abhisar):

Here weight is 600, so force require to lift it will be 60X10=600N

OpenStudy (abhisar):

Now work = 600X1.2=720J

OpenStudy (abhisar):

Power=720/1.5=480W

OpenStudy (abhisar):

LOL..damn typo..weight is 60Kg so Force = 60X10=600

OpenStudy (abhisar):

got it ?

OpenStudy (anonymous):

yes i do

OpenStudy (abhisar):

Sure ?

OpenStudy (anonymous):

yep

OpenStudy (abhisar):

That's gr8 ! \(\color{green}{\huge\ddot\smile}\)